### v1

# Single nuclear spin detection and control in a van der Waals material

Abstract
Optically active spin defects in solids [1, 2] are leading candidates for quantum sensing [3, 4] and quantum networking [5, 6]. Recently, single spin defects were discovered in hexagonal boron nitride (hBN) [7, 8, 9, 10, 11], a layered van der Waals (vdW) material. Due to its two-dimensional structure, hBN allows spin defects to be positioned closer to target samples than in three-dimensional crystals, making it ideal for atomic-scale quantum sensing [12], including nuclear magnetic resonance (NMR) of single molecules. However, the chemical structures of these defects [7, 8, 9, 10, 11] remain unknown, and detecting a single nuclear spin with an hBN spin defect has been elusive. In this study, we created single spin defects in hBN using 13C ion implantation and identified three distinct defect types. We observed both and spin states within a single hBN spin defect, with only the states showing strong hyperfine interactions with nearby 13C nuclear spins. For the first time, we demonstrated atomic-scale NMR and coherent control of individual nuclear spins in a vdW material. By comparing experimental results with density-functional theory calculations, we propose chemical structures for these spin defects. Our work advances the understanding of single spin defects in hBN and provides a pathway to enhance quantum sensing using hBN spin defects with nuclear spins as quantum memories.
Solid-state spin defects have become a leading platform for a wide range of quantum technologies, including multinode quantum networking [5, 6] and quantum-enhanced sensing [3, 4]. These advances are largely driven by the spin-photon quantum interfaces, which utilize optically addressable coherent spins. Despite the success of various spin-photon systems, each material platform has intrinsic limitations, creating trade-offs depending on the specific application [1, 2]. Moreover, spin defects that operate at room temperature are rare. Recently, optically active spin defects in hexagonal boron nitride (hBN), a van der Waals (vdW) material, has garnered vast attention [13, 7, 8, 9, 10, 11]. The layered structure of hBN facilitates integration with nanophotonic devices [14] and provides an ideal platform for quantum sensing at the atomic scale [12]. hBN spin defects have been used for sensing magnetic fields [15, 16, 17], temperature [16, 15], strain [18], and beyond [19]. However, electron spins in hBN suffer from short spin coherence times [20, 21, 22].
Nuclear spins typically exhibit long coherence times thanks to their weak coupling with the local environment, making them ideal candidates for quantum registers [23]. By employing nuclear spins as ancillary qubits, we can overcome the limitation of electron spin coherence times and enhance the sensitivity of a spin-based quantum sensor [24, 25]. This approach requires the ability to initialize, control, and read out individual nuclear spins [26, 27]. In the context of hBN spin defects, previous experiments used negatively charged boron vacancy () defects to polarize and read out nitrogen nuclear spin ensembles [28, 29, 30, 31]. However, these experiments were limited to ensemble-level operations due to the low quantum efficiency of defects. Moreover, the relatively short relaxation time of electron spins ( 20 s at room temperature) restricted the maximum operational time of nuclear spins. Recently discovered single spin defects in hBN [7, 8, 9, 10, 11] enabled the readout of individual electron spins in hBN. However, the chemical structures of these single spin defects remain unidentified, and the control of individual nuclear spins in hBN or other vdW materials is still elusive.
In this article, we report the first realization of single nuclear spin detection and control using a carbon-related defect in hBN. Our study reveals three major types of spin defects in 13C implanted hBN, with remarkably high optically detected magnetic resonance (ODMR) contrasts of up to 200. Two of these defect types are strongly coupled to nearby 13C nuclear spins. We observe the coexistence of and states within a single spin defect, though only the states exhibit strong coupling to 13C nuclear spins, with coupling strengths reaching up to 300 MHz. The strong hyperfine coupling results in well resolved hyperfine structures, enabling the initialize and readout of single nuclear spins assisted by a defect electron spin. We also demonstrate robust coherent control of a 13C nuclear spin and perform Ramsey and Hahn Echo measurements of the nuclear spin. The dephasing and coherence times of a typical 13C nuclear spin in hBN at room temperature are measured to be T = 16.6 s and T2 = 162 s, respectively. Our density-functional theory calculations suggest that CC donor-acceptor pairs (DAP) and CBON are likely candidates for the two types of spin defects with large hyperfine interactions.
Single spin defects in 13C implanted hBN
We create carbon-related spin defects in hBN [8, 9, 11] (Fig. 1) by 13CO2 (99 13C) ion implantation and thermal annealing. A confocal photoluminescence (PL) map reveals isolated emitters, as depicted in Fig. 1(j). Their optical spectra spread from 570 nm to 700 nm, depending on different defects (Figure 1(d)-(f)). Photon correlation measurements suggest that some defects are single photon emitters (). To verify the spin properties, we perform ODMR measurements by turning the microwave on and off while collecting the emitted photons. The ODMR contrast is determined by the ratio of photon count rates when the microwave is on () or off (): . In contrast to previous reports of spin defects with only a single resonant peak in ODMR [7, 8, 9, 10], we observe three distinct ODMR spectra with multiple resonances (Fig. 1(g)-(i)). These ODMR spectra feature a center branch (shaded areas in Fig. 1(g)-(i)) comprising multiple resonances that are assigned to hyperfine structures, along with additional side resonances located approximately 1 GHz away from the center branch. To understand the origination of the multiple resonances shown in the ODMR spectra, we first describe the system using the spin Hamiltonian
where and are the electron spin and the nuclear spin gyromagnetic ratios, and is the external magnetic field. denotes the electronic spin operator with or . and together are the zero-field splitting (ZFS) parameters, which are non-zero for the states (Figure 1(b)) but vanish for the states (Figure 1(c)). denotes the nuclear spin operator for the -th nuclear spin coupled to the defect electron. Specifically, for 13C nuclear spins, for 14N nuclear spins, and for 11B nuclear spins.
Based on ODMR spectra as shown in Fig. 1(g)-(i), we categorize the spin defects into three major groups, Group I-III, according to the number of peaks in the center branch. Individual peaks are labeled based on the defect group. The hyperfine splitting of Group II and III defects are determined to be 130 MHz and 300 MHz, respectively. It is worth noting that the 300 MHz hyperfine splitting is determined by the separation between the transitions III-2 and III-4. As will be discussed later, we assign the center peak III-3 to a second electron spin that weakly couples to the primary electron spin, suggested by the recently proposed spin pair model [32, 33]. The ODMR contrast can be as high as 200. A good single hBN spin defect in our sample has a typical sensitivity of 5 T/ for DC magnetic field sensing, calculated by [29], where is linewidth (20 MHz), is the contrast (30 ) and is the photon count rate (170 kcts/s).
To better understand the spin transitions, we measure the ODMR spectra as functions of out-of-plane magnetic fields (Figure 1 (k) and (i)). The results reveal an absence of zero field splitting (ZFSs) in the center branch, suggesting a spin manifold. In contrast, the side resonances (II-1, II-4, III-1 and III-5) confirm non-zero ZFSs and disperse with a -factor of 2. These resonances can be described by the spin Hamiltonian of Eq. 1, with a finite (typically 1 GHz) and along an out-of-plane quantization axis. This is distinct from the recent observation about a spin-triplet defect with an in-plane quantization axis [11], making it easier for magnetic field alignment. The additional peak II-5 of Defect 2 locates at the frequency = and has an anti-level crossing with II-4 at 37 mT, implying a double quantum transition. The assignment of and spin manifolds for these transitions is further supported by the measured ratio of Rabi oscillation frequencies () when the microwave is applied at line II-1 () and II-2 ().
Detection of single 13C nuclear spins in hBN
The well-resolved hyperfine structures enable us to distinguish different defect groups through two-dimensional (2D) mapping of the ODMR contrast at specific microwave driving frequencies. Figure 2 (a)-(c) present the ODMR contrast distribution of a 13C-implanted hBN with microwaves resonant at I-2, II-2 and III-2 transitions (Fig. 1(g)-(i)), respectively. All three maps reveal multiple spots with finite contrasts, enabling us to locate Group II and Group III defects. In contrast, no ODMR signal is observed in 12C-implanted hBN when microwaves are applied at the II-2 and III-2 transitions.
Besides ODMR, optically detected nuclear magnetic resonance (ODNMR) (Figure 2(g)) of 13C nuclear spins can provide further insights into the electron-nuclear hyperfine coupling. Here we conduct these measurements at the center branch of Group II and III defects. For both Defect 2 and Defect 3, we observe two closely spaced resonances in ODNMR at approximately half the frequencies of the hyperfine splitting observed in ODMR (Figure 2(h)-(i)). Such a feature verifies the configuration of the defect electron spin for these transitions. For a configuration, we expect the ODNMR resonance frequency to be the same as the hyperfine splitting in ODMR. The two-peak structure originates from the external magnetic field and/or a weaker hyperfine coupling to a second electronic spin. We also notice that different nuclear resonance frequencies are obtained when the microwave is applied at III-2 (III-4) and III-3. These features further verify two electron spins are involved in Defect 3.
Coherent control of a 13C nuclear spin in hBN
The well-isolated electron spin transitions at III-2 and III-4 (Fig. 1(i)) enable coherent manipulation of individual electron spin states associated with different nuclear spin states. This capability allows for the initialization, control, and readout of a nuclear spin via the electron spin, without the need for strong magnetic fields or level anticrossing (LAC) [28, 29, 30, 31].
To polarize the 13C nuclear spin, we first use a laser pulse to initialize the electron state (assuming ). Next, we apply a SWAP gate (Figure 3(a)) to swap the electron spin state and nuclear spin state, after which the nuclear spin is initialized to (or ). To estimate the fidelity of nuclear spin initialization, we measure the electron spin ODMR signal after the SWAP gate and calculate the polarization according to the imbalance between III-2 and III-4 (Figure 3(b)). We determine the nuclear spin polarization to be approximately 43 (-33) using the equation , where represents the amplitude of resonance III-j. A higher nuclear spin polarization of 60 can be achieved with another defect.
For nuclear spin coherent control, we use the protocol depicted in Figure 3(c). After polarizing the nuclear spin, another 10-s laser pulse re-initializes the electron spin state. Subsequently, we park the frequency of the selective RF pulse while varying its pulse duration . Finally, the nuclear spin state is read out via the spin defect using a selective microwave pulse and a 5-s laser pulse. Figure 3(d) shows the resulting nuclear spin Rabi oscillations. The oscillation persists for 40 s without significant decay, yielding a -gate fidelity of 99.75%. The extended operational time for nuclear spins is attributed to the long electron spin relaxation time (T1,e = 144 s) of this carbon-related spin defect, which is an order of magnitude longer than that of spin ensembles [28, 29]. By repeating the measurements at different RF powers, we observe a clear power dependence of the oscillations (Figure 3(f)), where the nuclear Rabi frequency is linearly proportional to the amplitude of the RF field (Figure 3(g)).
We further characterize the nuclear spin coherence via Ramsey and Hahn Echo sequences, as illustrated in Figure 4 (a)-(b). In the Ramsey interferometry, when RF pulses are applied exactly in resonance with the nuclear spin transitions, we observe a slow decay in 20 s. With a slight detuning of the RF frequency, an additional oscillation is observed alongside the original decay. By fitting the oscillation using the equation cos, we determine the inhomogeneous dephasing time to be T = 16.6 s and the oscillation frequency to be the same as the detuning. Figure 4(d) shows the exponential decay measured by the nuclear spin Hahn Echo sequence, revealing a nuclear spin coherence time of T2,n = 162 s. The nuclear spin coherence time is comparable with the electron spin relaxation time T1,e = 144 s, suggesting the measurement is limited by the electronic spin lifetime.
Identifying chemical structures
To identify possible chemical structures of the observed single spin defects, we perform first-principles calculations for the static properties and hyperfine interaction tensors. These properties can be directly compared with experimental results. Recent studies suggested the metastable configuration of the spin manifolds, while the optical transitions are supposed to come from the singlet state () [32, 33]. Therefore, for the simulation of ODMR using hyperfine tensors, we mainly focus on the spin manifolds. And we compare the optical spectra to the calculated ZPLs in the singlet manifolds.
Among the calculated candidates, CBCN-DAP-L [34] is a likely candidate for Group II defects (Figure 5(a)). This defect has a CBCN dimer structure with a larger separation ( lattice distance) between the two point defects. The positively charged state (CC-DAP-) contains an manifold, with a predicted hyperfine coupling strength of 135 MHz for the 13CN nuclear spin, closely matching the observed hyperfine coupling of 130 MHz. The simulated ODMR spectrum also aligns well with the measured ODMR, as shown in Figure 5(b). Since the spin density is primarily localized at the CN site and is barely affected by the distance , the hyperfine coupling strength remains nearly uncharged for different values. In contrast, the zero-photon line (ZPL) of CC-DAP- () is strongly affected by the distance . For example, the ZPL ranges from 1.75 eV to 2.14 eV as varies from to 4, agreeing well with our observed peak wavelength ranging from 585 nm (2.12 eV) to 700 nm (1.77 eV).
The CBON defect can be the spin defect in the Group III defects that contributes to the large hyperfine splitting. The positively charged state () has a spin doublet manifold () (Fig. 5(c)) with a hyperfine interaction strength of Azz = 314 MHz, which is close to the measured hyperfine coupling of 300 MHz. In Fig. 5(d), the simulated ODMR linewidth induced by the nuclear spin bath is 23 MHz, which is also close to the measured linewidth (16 MHz) under a weak microwave drive. The center peak (III-3) in Fig. 5(d) is attributed to a nearby second spin defect.
Conclusion
In conclusion, we report the detection and coherent control of single 13C nuclear spins using single hBN spin defects at room temperature. We find three distinct defect groups in 13CO2 implanted hBN samples, categorized based on their ODMR spectra. We observe both and spin states within a single hBN spin defect, which displays quantum coherence at room temperature and ODMR contrast up to 200. In addition, the electronic spin state can be readout using a reasonably long laser pulse of around 5 s, yielding approximately one photon per readout pulse and a single-short spin readout fidelity of = 0.12 (see Methods).
Leveraging the control of individual resonances within the well-resolved hyperfine structures, we demonstrate initialization, coherent control and readout of a single 13C nuclear spin using spin defects in Group II and Group III. The nuclear spins exhibits coherence times that are orders of magnitudes longer than those of electronic spins in hBN, offering the potential for long-lived quantum registers. The well-resolved hyperfine structure, combined with the high readout efficiency of spin states enabled by the high ODMR contrasts and the extended nuclear spin coherence times, makes this approach promising for achieving single-shot readout of individual nuclear spins [26]. This capability is crucial for implementing quantum error correction protocols in a quantum register [27]. Moreover, the 13C nuclear spin can serve as a quantum memory to enhance quantum sensing with single hBN spin defects.
Methods
Sample preparation
The hBN thin flakes were tape-exfoliated from a monocrystalline hBN crystal and transferred onto Si/SiO2 substrates. Then we irradiated the hBN flakes with 2.5 keV 13CO2 (99.0 13C, Sigma-Aldrich) ions using a home-built ion implanter. The sample is then annealed at high temperature at 10-5 torr for 2 hours to activate the carbon-related defects. For ODMR measurements, we transferred the hBN flakes to a coplanar waveguide using the standard dry transfer method with propylene carbonate stamps. The waveguide is made of 200 nm thick silver with a 4 nm thick Al2O3 layer on top.
Estimation of nuclear spin polarization
We estimate the polarization of the 13C nuclear spin by evaluating the imbalance between III-2 and III-4 in the ODMR spectrum. The ODMR is taken after the SWAP gate to transfer the electron polarization to the 13C nuclear spin. By using the fitted relative populations of the hyperfine basis states, the polarization can be calculated by the equation
Spin readout fidelity
The fidelity of a single-shot spin readout is an important factor to estimate how efficiently one can determine the electronic spin state of a spin defect, which is highly dependent on the defect properties. The fidelity is defined by the signal-to-noise ratio from a single readout pulse and can be express as [35]
where and are the mean numbers of detected photons for a single measurement of the brighter state and darker state, respectively. We estimate the fidelity based on the pulsed ODMR measurements. The pulsed ODMR contrast is 18 when we set readout duration at 5 s under a PMW=60 mW microwave drive. The contrast reaches 30 when PMW=2 W. For each readout laser pulse, we obtain approximately 1 photon from the darker state under the 15 W laser pumping. These yield the fidelity of 0.08 and 0.12 for PMW=60 mW and PMW=2 W, respectively.
Density-functional theory calculations
We use Quantum Espresso (QE) [36], an open-source plane-wave software, to perform the density-function theory (DFT) calculation. Both the Perdew-Burke-Ernzerhof (PBE) functional and the Heyd-Scuseria-Ernzerhof (HSE) hybrid functional (the factor of 0.32 for Fock exchange) [37, 38] are employed for the exchange-correlation interaction. We use Optimized Norm-Conserving Vanderbilt (ONCV) pseudopotential [39, 40] for the calculations of excitation energy, and the GIPAW pseudopotential [41] for the calculation of hyperfine interaction parameters and zero-field splitting (ZFS). We set the kinetic energy cutoff to be 55 Ry, which is adequate for converging the relevant properties. Geometry optimizations are carried out with a force threshold of 0.001 RyBohr. We select the or higher supercell size of hBN for the calculations of hyperfine parameter and excitation energies. For these calculations, we sample a k-point mesh of for the calculation of excitation energies [42], and point for the hyperfine parameters and ZFS [28, 42]. We calculate the zero-phonon line (ZPL) by the constraint occupation DFT (CDFT) method [43], the hyperfine parameters using the QE-GIPAW code [44], the ZFS by using the ZFS code [42], and we cross compare results between ZFS code and the PyZFS code [45].
Simulation of ODMR spectrum
The continuous wave (cw) ODMR spectra are simulated using the MATLAB toolbox EASYSPIN [46] based on data from the Ab initio calculations. EASYSPIN also takes the nuclear Zeeman and quadrupole interaction into account. Therefore, the cw ODMR linewidth can be determined according to the hyperfine couplings with the most abundant nuclear-spin-active isotopes: 13C,11B and 14N. In our simulation, we consider a 13C nuclear spin, ten nearest 11B nuclear spins and two proximate 14N nuclear spins. The other nuclei, located further away, couple more weakly to the electron, scaling with (where is the distance from the central carbon site), and thereby have a negligible effect on the ODMR linewidth.