### v2

# Single nuclear spin detection and control in a van der Waals material

Abstract
Optically active spin defects in solids [1, 2] are leading candidates for quantum sensing [3, 4] and quantum networking [5, 6]. Recently, single spin defects were discovered in hexagonal boron nitride (hBN) [7, 8, 9, 10, 11], a layered van der Waals (vdW) material. Due to its two-dimensional structure, hBN allows spin defects to be positioned closer to target samples than in three-dimensional crystals, making it ideal for atomic-scale quantum sensing [12], including nuclear magnetic resonance (NMR) of single molecules. However, the chemical structures of these defects [7, 8, 9, 10, 11] remain unknown, and detecting a single nuclear spin with an hBN spin defect has been elusive. In this study, we created single spin defects in hBN using 13C ion implantation and identified three distinct defect types based on hyperfine interactions. We observed both and spin states within a single hBN spin defect. We demonstrated atomic-scale NMR and coherent control of individual nuclear spins in a vdW material, with a -gate fidelity up to 99.75% at room temperature. By comparing experimental results with density-functional theory calculations, we propose chemical structures for these spin defects. Our work advances the understanding of single spin defects in hBN and provides a pathway to enhance quantum sensing using hBN spin defects with nuclear spins as quantum memories.
Solid-state spin defects have become a leading platform for a wide range of quantum technologies, including multinode quantum networking [5, 6] and quantum-enhanced sensing [3, 4]. These advances are largely driven by the spin-photon quantum interfaces, which utilize optically addressable coherent spins. Despite the success of various spin-photon systems, each material platform has intrinsic limitations, creating trade-offs depending on the specific application [1, 2]. Moreover, spin defects that operate at room temperature are rare. Recently, optically active spin defects in hexagonal boron nitride (hBN), a van der Waals (vdW) material, has garnered vast attention [13, 7, 8, 9, 10, 11]. The layered structure of hBN facilitates integration with nanophotonic devices [14] and provides an ideal platform for quantum sensing at the atomic scale [12]. hBN spin defects have been used for sensing magnetic fields [15, 16, 17], temperature [16, 15], strain [18], and beyond [19]. However, electron spins in hBN suffer from short spin coherence times [20, 21, 22].
Nuclear spins typically exhibit long coherence times thanks to their weak coupling with the local environment, making them ideal candidates for quantum registers [23]. By employing nuclear spins as ancillary qubits, we can overcome the limitation of electron spin coherence times and enhance the sensitivity of a spin-based quantum sensor [24, 25]. This approach requires the ability to initialize, control, and read out individual nuclear spins [26, 27]. In the context of hBN spin defects, previous experiments used negatively charged boron vacancy () defects to polarize and read out nitrogen nuclear spin ensembles [28, 29, 30, 31]. However, these experiments were limited to ensemble-level operations due to the low quantum efficiency of defects. Moreover, the relatively short relaxation time of electron spins ( 20 s at room temperature) restricted the maximum operational time of nuclear spins. Recently discovered single spin defects in hBN [7, 8, 9, 10, 11] enabled the readout of individual electron spins in hBN. However, the chemical structures of these single spin defects remain unidentified, and the control of individual nuclear spins in hBN or other vdW materials is still elusive.
In this article, we report the first realization of single nuclear spin detection and control using a carbon-related defect in hBN. Our study reveals three major types of single spin defects in 13C implanted hBN, with remarkably high optically detected magnetic resonance (ODMR) contrasts of up to 200. Two of these defect types are strongly coupled to nearby 13C nuclear spins. We observe the coexistence of and states within a single spin defect, though only the states exhibit strong coupling to 13C nuclear spins, with coupling strengths reaching up to 300 MHz. The strong hyperfine coupling results in well resolved hyperfine structures, enabling the initialization and readout of single nuclear spins assisted by a defect electron spin. We realize electron-nuclear spin two-qubit gate and coherent control of a 13C nuclear spin, with a nuclear spin -gate fidelity up tp 99.75% at room temperature. We also perform Ramsey and Hahn Echo measurements of the nuclear spin. The dephasing and coherence times of a typical 13C nuclear spin in hBN at room temperature are measured to be = 16.6 s and = 162 s, respectively. Our density-functional theory calculations suggest that CC donor-acceptor pairs (DAP) and CBON are likely key components for the two types of spin defects with large hyperfine interactions.
Single spin defects in 13C implanted hBN
We create carbon-related spin defects in hBN [8, 9, 11] (Fig. 1) by 13CO2 (99 13C) ion implantation and thermal annealing (Supplemental Section II). A confocal photoluminescence (PL) map reveals isolated emitters, as depicted in Fig. 1(b). Their optical spectra spread from 570 nm to 700 nm, depending on different defects (Figure 1(d)-(f)). Photon correlation measurements suggest that some defects are single photon emitters (). To verify their spin properties, we perform ODMR measurements by toggling the microwave on and off while collecting the emitted photons. The ODMR contrast is determined by the ratio of photon count rates when the microwave is on () or off (): . In contrast to previous reports of spin defects with only a single resonant peak in ODMR [7, 8, 9, 10], we observe three distinct ODMR spectra with multiple resonances (Fig. 1(g)-(i)). The positive contrast indicates that the defects are initialized into a darker state, while we also observe defects with negative contrasts but are rarer (Supplemental Figure S20). These ODMR spectra feature a center branch (shaded areas in Fig. 1(g)-(i)) comprising multiple resonances that are assigned to hyperfine structures, along with additional side resonances located approximately 1 GHz away from the center branch. The high ODMR contrast and brightness yield a typical DC magnetic field sensitivity of 5 T/ (Methods).
Based on ODMR spectra as shown in Fig. 1(g)-(i), we categorize the spin defects into three major groups, Group I-III, according to the number of peaks in the center branch. Individual peaks are labeled according to their respective defect group. The hyperfine splitting of Group II and III defects are determined to be 130 MHz and 300 MHz, respectively. Notably, the 300 MHz hyperfine splitting is determined by the separation between transitions III-2 and III-4. The additional center peak, III-3, exhibits a different hyperfine stucture under weak microwave driving, suggesting that it originates from a second electron spin. This observation supports the recently proposed spin-pair model, in which a single defect complex hosts two electron spins, separated by several nanometers, that couple differently to the 13C nuclear spins (Fig. 1(j)) [32, 33].
To better understand the spin transitions, we measure the ODMR spectra as a function of out-of-plane magnetic fields (Figure 1(k) and 1(i)). The results reveal an absence of zero-field splitting (ZFS) in the center branch, corresponding to the transition within the spin manifold (see Supplemental Section IV for experimental verification). In contrast, the side resonances (II-1, II-4, III-1 and III-5) exhibit non-zero ZFS and disperse with a -factor of 2, confirming that they correspond to the transitions within the spin manifold. The Hamiltonian of each spin manifold is given by:
where and are the electron spin and the nuclear spin gyromagnetic ratios, and is the external magnetic field. denotes the electronic spin operator with or . and together are the ZFS parameters, which are non-zero for the states but vanish for the states. denotes the nuclear spin operator for the -th nuclear spin coupled to the defect electron. Specifically, for 13C nuclear spins, for 14N nuclear spins, and for 11B nuclear spins. The hyperfine interaction between the electron spin and the i-th nuclear spin is described by the interaction tensor .
The magnetic field dependent ODMR allows us to extract the ZFS parameter (typically 1 GHz) and (varies from 100 MHz to 400 MHz) along an out-of-plane quantization axis for the S = 1 transitions (Supplementary Figure S3). This is different from the recent observation of a spin-triplet defect with an in-plane quantization axis [11], making our system more favorable for magnetic field alignment. Additionally, an extra peak II-5 of Defect 2 locates at the frequency = and has an LAC with II-4 at 37 mT (Supplemental Figure S3), suggesting a double-quantum transition. While previously proposed spin-pair models based on DAP have focused solely on S = 1/2 transitions in metastable state (MS) [32, 33], the coexistence of both S = 1/2 and S = 1 transitions can be well explained by an extended spin-pair model incorporating additional energy levels and transitions (see Methods). The DAP model also accounts for the large variation in PL spectra observed across different defects, despite their similar ODMR signatures [34, 35].
Detection of single 13C nuclear spins in hBN
The well-resolved hyperfine structures enable us to distinguish different defect groups through two-dimensional (2D) mapping of the ODMR contrast at specific microwave driving frequencies. Figure 2 (a)-(c) present the ODMR contrast distribution of a 13C-implanted hBN with microwaves resonant at I-2, II-2 and III-2 transitions (Fig. 1(g)-(i)), respectively. All three maps reveal multiple spots with finite contrasts, enabling us to locate Group II and Group III defects. In contrast, no ODMR signal is observed in 12C-implanted hBN when microwaves are applied at the II-2 and III-2 transitions.
Besides ODMR, optically detected nuclear magnetic resonance (ODNMR) (Figure 2(g)) of 13C nuclear spins can provide further insights into the electron-nuclear hyperfine coupling. Here we conduct these measurements at the center branch of Group II and III defects. For both Defect 2 and Defect 3, we observe two closely spaced resonances in ODNMR at approximately half the frequencies of the hyperfine splitting observed in ODMR (Figure 2(h)-(i)). Such a feature verifies the configuration of the defect electron spin for these transitions. For a configuration, we expect the ODNMR resonance frequency to be the same as the hyperfine splitting in ODMR. The two-peak structure originates from the external magnetic field and/or a weaker hyperfine coupling to a second electronic spin. We also notice that different nuclear resonance frequencies are obtained when the microwave is applied at III-2 (III-4) and III-3 (Supplemental Figure S17). These features further verify two electron spins are involved in Defect 3.
Coherent control of a 13C nuclear spin in hBN
The well-isolated electron spin transitions at III-2 and III-4 (Fig. 1(i)) enable coherent manipulation of individual electron spin states associated with different nuclear spin states. This capability allows for the initialization, control, and readout of a nuclear spin via the electron spin, without the need for strong magnetic fields or level anticrossing (LAC) [28, 29, 30, 31].
To polarize the 13C nuclear spin, we first use a laser pulse to initialize the electron state (assuming ). Next, we apply a SWAP gate (Figure 3(a)) to swap the electron spin state and nuclear spin state, after which the nuclear spin is initialized to (or ). To estimate the fidelity of nuclear spin initialization, we measure the electron spin ODMR signal after the SWAP gate and calculate the polarization according to the imbalance between III-2 and III-4 (Figure 3(b)). We determine the nuclear spin polarization to be approximately 43 (-33) using the equation , where represents the amplitude of resonance III-j. A higher nuclear spin polarization of 60 can be achieved with another defect as shown in Supplemental Figure S18.
For nuclear spin coherent control, we use the protocol depicted in Figure 3(c). After polarizing the nuclear spin, another 10-s laser pulse re-initializes the electron spin state. Subsequently, we park the frequency of the selective RF pulse while varying its pulse duration . Finally, the nuclear spin state is read out via the spin defect using a selective microwave pulse and a 5-s laser pulse. Figure 3(d) shows the resulting nuclear spin Rabi oscillations. We determine a -pulse time of 0.6 s and a coherence time of 117 by fitting the oscillation contrast to , yielding a -gate fidelity of 99.75% (Methods). The extended operational time for nuclear spins is attributed to the long electron spin relaxation time ( = 144 s, Supplemental Figure S15) of this carbon-related spin defect, which is an order of magnitude longer than that of spin ensembles [28, 29]. By repeating the measurements at different RF powers, we observe a clear power dependence of the oscillations (Figure 3(f)), where the nuclear Rabi frequency is linearly proportional to the amplitude of the RF field (Figure 3(e)).
We further characterize the nuclear spin coherence via Ramsey and Hahn Echo sequences, as illustrated in Figure 4 (a)-(b). In the Ramsey interferometry, when RF pulses are applied exactly in resonance with the nuclear spin transitions, we observe a slow decay in 20 s. With a slight detuning of the RF frequency, an additional oscillation is observed alongside the original decay. By fitting the oscillation using the equation cos, we determine the inhomogeneous dephasing time to be = 16.6 s and the oscillation frequency to be the same as the detuning. Figure 4(d) shows the exponential decay measured by the nuclear spin Hahn Echo sequence, revealing a nuclear spin coherence time of = 162 s. The nuclear spin coherence time is comparable with the electron spin relaxation time = 144 s (Supplemental Figure S15), suggesting the measurement is limited by the electronic spin lifetime.
Identifying chemical structures
In addition to providing long-lived spin qubits, nuclear spin magnetic resonance offers valuable insight into identifying defect structures. Here, we perform first-principles calculations of the static properties and hyperfine interaction tensors to enable direct comparison with experimental data. Our calculations suggest that CBCN-DAP- [36] is a likely member of the spin pair responsible for Group II defects (Figure 5(a)). This dimer can exist in two charge states: CC-DAP- and CC-DAP-. The positively charged state (CC-DAP-) has an manifold, and our calculations predict a 13CN hyperfine coupling of 135 MHz (for ). This value closely matches the experimentally observed 130 MHz. The simulated ODMR spectrum aligns well with the measured ODMR, as shown in Figure 5(b). In addition, the calculated hyperfine coupling for the nearest 11B nuclear spins is 17.6 MHz, resulting in an ODNMR resonance at 8.8 MHz, which closely matches the observed ODNMR resonance at 9.2 MHz in Figure 2(h). Since the spin density is primarily localized at the CN site and is barely affected by the distance , the hyperfine coupling strength remains nearly uncharged for different values (Supplemental Figure S32).
To account for the observed 300 MHz hyperfine splitting in Group III defects, we consider an oxygen substitution next to the 13CB site, forming a CBON defect [10] as one member of the spin pair. Oxygen impurities exist in hBN samples [37] and can also be introduced during annealing due to residual oxygen in the vacuum chamber. The positively charged state (CBO) of this defect hosts a spin-doublet manifold () (Figure 5(c)) with a hyperfine interaction strength MHz, in close agreement with our measured 300 MHz splitting. In Figure 5(d), the simulated ODMR linewidth induced by the nuclear spin bath is 23 MHz, comparable to the 16 MHz linewidth measured under weak microwave driving (Supplemental Figure S13). The central peak (III-3) in Fig. 5(d)) is attributed to a second nearby spin defect.
For both Group II and III defects, the structures discussed above correspond to the defect members in the spin pairs that exhibit strong hyperfine coupling. We expect a second, nearby donor or acceptor defect, couples to the proposed defect candidates to form a defect DAP complex. The structure of the other defect remains unclear due to lack of characteristic hyperfine features. Further identification of both defect constituents may be possible by using high-purity hBN to suppress intrinsic 12C impurities, combined with higher 13C implantation density. This approach would improve the probability that both defects interact with 13C nuclear spins, enabling more definitive structural assignments.
Conclusion
In conclusion, we report the detection and coherent control of single 13C nuclear spins using single hBN spin defects at room temperature. We find three distinct defect groups in 13CO2 implanted hBN samples, categorized based on their ODMR spectra. We observe both and spin states within a single hBN spin defect complex, which displays quantum coherence at room temperature and ODMR contrast up to 200. In addition, the electronic spin state can be readout using a reasonably long laser pulse of around 5 s, yielding approximately one photon per readout pulse (Supplemental Figure S16) and a single-short spin readout efficiency of = 0.12 (see Methods).
Leveraging the control of individual resonances within the well-resolved hyperfine structures, we demonstrate initialization, coherent control and readout of a single 13C nuclear spin using spin defects in Group II (Supplemental Section IV) and Group III. The nuclear spins exhibits coherence times that are orders of magnitudes longer than those of electronic spins in hBN, offering the potential for long-lived quantum registers. The well-resolved hyperfine structure, combined with the high readout efficiency of spin states enabled by the high ODMR contrasts and the extended nuclear spin coherence times, makes this approach promising for achieving single-shot readout of individual nuclear spins [26]. This capability is crucial for implementing quantum error correction protocols in a quantum register [27]. Moreover, the 13C nuclear spin can serve as a quantum memory to enhance quantum sensing with single hBN spin defects.
ACKNOWLEDGMENTS
T.L. acknowledges the support by the Gordon and Betty Moore Foundation, grant DOI 10.37807/gbmf12259, and the National Science Foundation grant PHY-2409607. The investigation of the spin pair model was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences through QuPIDC Energy Frontier Research Center under award Number DE-SC0025620. Y.P. acknowledges the support by the National Science Foundation under grant no. DMR2143233. The ab initio calculations used resources of the Scientific Data and Computing Center, a component of the Computational Science Initiative, at Brookhaven National Laboratory under Contract No. DE-SC0012704, the Lux Supercomputer at UC Santa Cruz, funded by NSF MRI Grant No. AST 1828315, the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02- 05CH11231, and the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation Grant No. ACI-1548562.
AUTHOR CONTRIBUTIONS
T.L. and X.G. conceived and designed the project. X.G., S.V., K.S. and Y.J. built the setup. K.L. and S.Z. performed the ab initio calculations. P.J. fabricated microwave waveguides. S.V. and S.D. performed the ion implantation and thermal annealing on the hBN samples. X.G. and S.V. performed measurements. X.G. and Z.G. performed the numerical simulation. X.G., S.V., T.L., K.L. and Y.P. analyzed the results. T.L. and Y.P. supervised the project. All authors contributed to the writing of the manuscript.